Home‎ > ‎Crew 161‎ > ‎

0120-DailySummaryReport

MDRS Daily Summary Report

Sol 11 1/20/2016
 
Martian geology.
 
Executive Officer Reporting
 
Mission Status: Research projects are on track and mission is going on smoothly.
 
Sol Activity Summary: Today, we tested our robot again on the Martian landscape. Specifically, the scooping function of the robot was tested. The engineering team is not completely satisfied with results and will be conducting a further test tomorrow,
 
Look Ahead Plan: An EVA is planned for tomorrow.
 
Anomalies in work:  One of the ATVs broke down during usage today. Crew engineer has a more detailed information on the situation.
 
Weather: The Martian weather was warm, with clear skies.
 
Crew Physical Status:  All crew members are hale and hearty! Crew HSO report details current status of all crew members.
 
EVA: EVA activity was successful. Soil around Wash was collected for further studies.
 
Reports to be filed:
Commanders Report
EVA #9 report
Daily Summary Report
Journalist Report
Biologist Report
Engineering Report
EVA #10 request

Support Requested: Engineering support is requested for ATV.

Best regards,

Adeniyi Adebayo
XO, Crew 161